OpenStudy (anonymous):

Solve the equation 2(x^4)+3(x^3)-7(x^2)+3x-9=0 by first finding any rational roots. Thanks!

OpenStudy (anonymous):

possible rational roots are \[\pm1, \pm3, \pm9, \pm\frac{1}{2}, \pm\frac{3}{2},\pm\frac{9}{2}\]

OpenStudy (anonymous):

a lot to check actually so i would start with the easy ones : 1 and -1

OpenStudy (anonymous):

1 does not work

OpenStudy (anonymous):

if you give me a minute i will find them for you

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

-1 does not work either

OpenStudy (anonymous):

don't you need to use synthetic division to find the other roots as well?

OpenStudy (anonymous):

well you need to find at least one to start. you can use synthetic division sure. i was going to cheat

OpenStudy (anonymous):

-3 works

OpenStudy (anonymous):

once you have -3 you can use synthetic division to factor

OpenStudy (anonymous):

and then what do I do w/ the depressed equation?

OpenStudy (anonymous):

depressed?

OpenStudy (anonymous):

you will get \[p(x)=(x+3)q(x)\]

OpenStudy (anonymous):

you find \[q(x)\] by dividing

OpenStudy (anonymous):

in fact i have done it.

OpenStudy (anonymous):

would you like the answer or do you want to find it?

OpenStudy (anonymous):

both. if you do not mind.

OpenStudy (anonymous):

ok. first of all \[p(-3)=0\] you can check this by synthetic division, or by replacing x by -3 which is kind of a pain

OpenStudy (anonymous):

ok, i have done the division. the equation is 2x^3-3x^2+2x-3=0.

OpenStudy (anonymous):

k you were ahead of me

OpenStudy (anonymous):

now the possible rational roots have narrowed somewhat. we know 1 and -1 don't work. we found -3. 3 does not work. now you can try \[\frac{3}{2}\] and \[-\frac{3}{2}\]

OpenStudy (anonymous):

i give you a hint, it is not the negative one!

OpenStudy (anonymous):

try 3/2 and you will get no remainder

OpenStudy (anonymous):

yes, i found that.

OpenStudy (anonymous):

when do I stop finding roots?

OpenStudy (anonymous):

ok and you get \[p(x)=(x+3)(x-\frac{3}{2})(2x^2+2)\]

OpenStudy (anonymous):

or \[p(x)=(x+3)(2x-3)(x^2+1)\]

OpenStudy (anonymous):

now you can stop because \[x^2+1\] does not factor over the real numbers

OpenStudy (anonymous):

if you are using complex numbers then i guess you can write \[p(x)=(x+3)(2x-3)(x+i)(x-i)\] but that is kind of silly because we are working with a real valued polynomial

OpenStudy (anonymous):

is it also b/c the original equation is to the third degree? somehow, I also found +i, -i...

OpenStudy (anonymous):

yes because the roots of \[x^2+1\] are i and -i

OpenStudy (anonymous):

we have found the 4 zeros, 2 real and two complex. of course you can stop now because a polynomial of degree 4 can have a most 4 zeros

OpenStudy (anonymous):

ohhhh, so +i and -i can b considered any number?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

they have a specific meaning

OpenStudy (anonymous):

square root of -1.

OpenStudy (anonymous):

\[x^2+1=0\] \[x^2=-1\] \[x=\pm\sqrt{-1}\]

OpenStudy (anonymous):

it is just a symbol for \[\sqrt{-1}\]

OpenStudy (anonymous):

ahhhhhh. ok. thanks a ton!!

OpenStudy (anonymous):

not a real number

OpenStudy (anonymous):

welcome. hope it was clear

OpenStudy (anonymous):

definitely.